The point on the hyperbola x2/24-y2/18=1 which is nearest to the line 3x+2y+1= 0 is

Grade:11

2 Answers

Khimraj
7 years ago
The nearest point will be that at which tangent on hyperbola is parallel to given line. Slope of parabola is y` = 3x/4y = -3/2 then x = -2y then put this equation of parabola and solve then you find coordinate (x,y) = (4√30,6√2). Hope it clears. If still u have doubts then please clarify. If u like answer then please approve the answer.
Khimraj
7 years ago
Sorry for correction by putting x= -2y and solve u get coordinate (x,y) =(-6,3). Hope now it clears. If still u have doubts then please clarify. If u like answer then please approve the answer.